Tthe Winston Company esmates that the tactery overhead for the following year wit be 5346,400 The company has decided that the basis for noolyieng factory overhead should be machine hours, which is estimatod to be 36,800 hours. The total machine hours for the year were 54,900. The actual fectory overhead for the year wre $1,255.000. tho totxl ชctory overhead anourit appled. (Round to the nearest dalac) h Caloate the over or underapoled anount for the yeat Enter the amount an a positive number Prepare the jourrol entry to dese factory overhead inte Cost of Goods Seld
Answer #1
a) Predetermined OH rate = 846400/36800 = $            23.00
Total factory overhead applied = 54900*23 = $    12,62,700
b) Overapplied overhead = 1262700-1255000 = $            7,700
c) Factory overhead $            7,700
Cost of goods sold $       7,700